Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

A probablity question...(its easy but i am not able 2 solve)

Expert replies
by mathurmudit1003 » Wed Jun 20, 2012 1:08 pm
In a bag of marbles, there are 3 red, 2 white, and 5 blue. If Bob takes 2 marbles out of
the bag, what is the probability that he will have one white and one blue marble?

need explanation for the solution
Join the discussion
Source: — Problem Solving |

by eagleeye » Wed Jun 20, 2012 3:02 pm
Hi mathurmudit1003:

There are 2 ways in which you can easily do it:
1. Selections:

Here we have 3R, 2W, 5R for a total of 10 marbles. We need to find the probability of selecting one white and one blue.

a. No. of ways of selecting 2 marbles out of 10 = 10C2 = 45
b. No. of ways of selecting 1 white marble out of 2 = 2C1 = 2
c. No. of ways of selecting 1 red marble out of 5 5C1 = 5

Now the probability of selecting one white AND one red = (2C1*5C1)//(10C2) = 2*5/45 = 2/9.

2. Arrangements:
We need to select two marbles – one white and one blue. There are two arrangements (order is important) that you can do with those.

a. Select white one first, select red one second.
b. Select red one first, select white one second.

Now probability of selecting the first one white = (No. of whites)/(Total) = 2/10
Probability of selecting the second one red (after selecting the first one) = (No. of Reds)/(Remaining Total) = 5/9.

So probability of selecting first white AND second red = 2/10*5/9.
Same way probability of selecting red one first AND white one second = 5/10*2/9.

So probability required = SUM of the two cases above = 2/10*5/9 + 5/10*2/9 = 20/90 = 2/9.
Let me know if this helps :)
Join the discussion